# Gradient Guidance for Diffusion Models: An Optimization Perspective

### v1

@@ Abstract
Diffusion models have demonstrated empirical successes in various applications and can be adapted to task-specific needs via guidance. This paper introduces a form of gradient guidance for adapting or fine-tuning diffusion models towards user-specified optimization objectives. We study the theoretic aspects of a guided score-based sampling process, linking the gradient-guided diffusion model to first-order optimization. We show that adding gradient guidance to the sampling process of a pre-trained diffusion model is essentially equivalent to solving a regularized optimization problem, where the regularization term acts as a prior determined by the pre-training data. Diffusion models are able to learn data’s latent subspace, however, explicitly adding the gradient of an external objective function to the sample process would jeopardize the structure in generated samples. To remedy this issue, we consider a modified form of gradient guidance based on a forward prediction loss, which leverages the pre-trained score function to preserve the latent structure in generated samples. We further consider an iteratively fine-tuned version of gradient-guided diffusion where one can query gradients at newly generated data points and update the score network using new samples. This process mimics a first-order optimization iteration in expectation, for which we proved $\mathcal{O}(1/K)$ convergence rate to the global optimum when the objective function is concave.
@@ 1 Introduction
Diffusion models have emerged as a significant advancement in the field of generative artificial intelligence, offering state-of-the-art performance in image generation (Song and Ermon, 2019; Song et al., 2020a; Dhariwal and Nichol, 2021). These models operate by gradually transforming a random noise distribution into a structured output, by using a score function trained from large amounts of data. Such a transforming process is typically modeled as a stochastic differential equation, offering a mathematically grounded approach for sampling. One of the key advantages of diffusion models is their ability to be guided or fine-tuned for specific tasks, which allows them to excel in a wide range of applications (Kong et al., 2020; Ajay et al., 2022; Gruver et al., 2023).
Guidance-based diffusion, a nuanced extension of diffusion models, stands at the forefront of controlled generation in generative AI. This approach involves steering the noise transformation process of a diffusion model towards desired outcomes by incorporating additional “guidance signals”. This guidance can manifest in various forms, such as text prompts, class labels, or even conditioning on specific attributes and rewards. The core principle behind this technique is to influence the probabilistic pathway of the noise transformation process at each time step, thereby steering the final output towards predefined criteria or objectives. This controlled generation capability opens up opportunities for generative AI in a broad range of tasks, such as in targeted image synthesis, content creation with specific themes, or even in drug design where molecular structures need to meet specifications.
A notable example is the classifier-based diffusion model introduced by Song et al. (2020c); Dhariwal and Nichol (2021), which generates data conditioned on a class label, via guidance signals that are computed from conditional likelihoods from a classifier. Building on this concept, Bansal et al. (2023) extend the classifier-guidance method to a form of “universal guidance”. Such guidance allows the generation process to be influenced by gradient obtained from some external loss function, effectively tailoring the diffusion process to meet specific objectives (Chung et al., 2022a, b; Graikos et al., 2022; Kawar et al., 2022; Lugmayr et al., 2022; Wang et al., 2022). Despite of numerous empirical successes, there remain significant gaps in the theoretical understanding and guarantees associated with guided diffusion models.
@@
Suppose we have a pre-trained diffusion model that can generate new samples faithfully from the pre-training data’s distribution and maintain the data’s latent structure. The goal is to adapt this diffusion model to generate new samples that optimize task-specific objectives, while maintaining the learned structure in new samples. Compared to classic optimization, the guided diffusion model offers new possibilities to optimize complex design variables such as images, videos, proteins, and genomes (Black et al., 2023; Watson et al., 2023; Liu et al., 2024). Interested readers may refer to recent surveys for a more comprehensive exposure (Yang et al., 2023; Chen et al., 2024; Guo et al., 2023).
To adapt pre-trained diffusion models, existing practical methods largely rely on empirical heuristics and hyperparameter tuning. There remain critical theoretical questions: (i) Why does naively guiding diffusion models using gradient never work in practice? (ii) How to add a guidance signal to improve the target objective without compromising the quality of the generated output? (iii) Can one guarantee the properties of new samples generated by guided diffusion? (iv) What are the limits of adaptability in these guided models? This paper aims to answer these questions from an optimization perspective.
@@
Scope of This Paper
We investigate the role of guidance in diffusion models from an optimization perspective. The goal is to generate samples that optimize a given objective function $f$. Drawing inspiration from gradient-based optimization methods, we construct a guidance signal based on the gradient vector, $\nabla f$. Then we use the gradient signal, in addition to the pre-trained score function, to guide the sampling process towards generating structured output with higher function values. See Figure 1 for illustration our algorithmic framework. Our main results are summarized as follows:
  • We focus on structured data. Assume that pre-trained data belongs to a latent low-dimensional subspace, thus the trained score function is capable of discerning and maintaining the latent subspace structure of data during (unguided) generation. To guide the generation, we introduce a gradient-like guidance based on a forward prediction loss (Definition 1). This gradient guidance can be computed based on a pre-trained score network, and it provably preserve any learned low-dimensional structure in the generated output (Lemma 1).
  • We formalize a mathematical framework of using a gradient-guided diffusion model for generative optimization. While retaining the pre-trained score function, the generation process is iteratively refined and guided using new gradient queries (Algorithm 1; Figure 1 without fine-tuning). Under proper assumptions, we demonstrate that this adapted model generates novel samples whose expectation converges to a solution that is regularized with respect to the original problem (Theorem 2). The regularization ensures that the generated samples remain proximal to the training data. In other words, gradient guidance cannot shift data distribution unboundedly towards higher objective values, revealing a fundamental limit for adapting pre-trained diffusion models.
@@
  • Furthermore, we explore an adaptive variant of gradient-guided diffusion, where both the score function and gradient guidance are iteratively fine-tuned using self-generated samples (Algorithm 2; Figure 1 with fine-tuning). Although slightly increasing the computational demand, we provide evidence that this approach generates new samples whose expectation converges to global optima, within the latent subspace, at a rate of $\mathcal{O}(1/K)$ (Theorems 3 and 4), where $K$ denotes the number of iterations and gradient evaluations, matching classical convergence theory of convex optimization.
Our findings suggest that this novel gradient guidance not only preserves the latent subspace structure of the data but also ensures fast convergence towards the optimal solution. Numerical experiments with a pre-trained U-network score function are provided in Section 7 to support these theoretical findings.
2 Related Work
Our study is motivated by recent empirical progress of guidance-based diffusion model fine-tuning for steering sample generation towards specific needs (Dhariwal and Nichol, 2021; Bansal et al., 2023). Upon modeling the specific needs as a reward function, the relevant methods can be summarized into two categories in the sequel.
Guided Generation and Fine-tuning
Given an auxiliary reward function judging the sample property of interests, existing research explored diverse mechanisms to guide generation from diffusion models. There are mainly two types of methods. The first type of method incorporates an additive so-called “guidance” term into the score function of pre-trained diffusion models at inference time. For example, classifier guidance (Song et al., 2020a; Dhariwal and Nichol, 2021) defines the guidance term as the gradient of an externally trained classifier on noise corrupted data. Classifier-free guidance (Ho and Salimans, 2022) simultaneously trains conditional and unconditional diffusion models, circumventing the training of an external classifier. After the training, the score functions of the conditional and unconditional models are combined together to achieve guided generation. Bansal et al. (2023) draw motivation from classifier guidance and generalize the idea to a “universal guidance” for adapting unconditioned score functions to various external rewards.
The second type of method attempts to directly fine-tune the weight parameters in a pre-trained diffusion model by interacting with the target reward function. For example, Clark et al. (2023) fine-tune diffusion models by directly backpropagating the gradient of the reward function. Recently, a line of works utilizes Reinforcement Learning (RL) techniques for fine-tuning diffusion models (Black et al., 2023; Fan et al., 2023). They formulate the sample generation process of diffusion models as a finite-horizon Markov chain. The score function can be viewed as a policy, the generated samples are the state of the Markov chain, and the target reward function defines the terminal reward. In this way, fine-tuning diffusion models is equivalent to policy optimization and allows the use of policy gradient methods.
Sampling and Statistical Theory of Diffusion Model
In contrast to the fruitful empirical advances, the theory of diffusion models is still limited. To the best of our knowledge, a theoretical understanding of fine-tuning diffusion models is absent. Existing results mainly focus on the sampling ability and statistical properties of unconditional diffusion models. In particular, for sampling ability, a line of works shows that the distribution generated by a diffusion model is close to the data distribution, as long as the score function is accurately estimated (De Bortoli et al., 2021; Albergo et al., 2023; Block et al., 2020; Lee et al., 2022a; Chen et al., 2022; Lee et al., 2022b, a; Chen et al., 2022; Lee et al., 2022b). The accuracy of the estimated score function is measured in terms of an $L_{\infty}$ or $L_{2}$-norm distance. More recently, Chen et al. (2023c, b); Benton et al. (2023) develop refined and tighter analyses using Taylor expansions of the discretized backward process and localization method. It is worth mentioning that the analysis in Chen et al. (2023c, b); Benton et al. (2023) extends to broader sample generation processes such as deterministic ones based on probabilistic ODEs. Going beyond distributions in Euclidean spaces, De Bortoli (2022) analyzes diffusion models for sampling distribution supported on a low-dimensional manifold. Moreover, Montanari and Wu (2023) consider sampling from symmetric spiked models, and El Alaoui et al. (2023) study sampling from Gibbs distributions using diffusion processes.
Turning towards the statistical theory of diffusion models, Song et al. (2020b) and Liu et al. (2022) provide asymptotic analyses, assuming a parametric form of the score function. Unfortunately, asymptotic analysis does not lead to concrete sample complexities. Later, concurrent works, Oko et al. (2023) and Chen et al. (2023a), establish sample complexity bounds of diffusion models for estimating nonparametric data distributions. In high dimensions, their results highlight a curse of dimensionality issue without further assumptions, which also appears in Wibisono et al. (2024) considering kernel methods. More interestingly, these works demonstrate that diffusion models can circumvent the curse of dimensionality issue if the data has low-dimensional structures. In the same spirit, Mei and Wu (2023) investigate learning high-dimensional graphical models using diffusion models, without the curse of dimensionality. For conditional diffusion models, Yuan et al. (2023); Fu et al. (2024) establish sample complexity bounds for learning generic conditional distributions. We refer readers to Chen et al. (2024) for an overview of contemporary theoretical progress.
Novelty of This Paper
Despite the existing theoretical underpinnings of diffusion models, our paper provides the first rigorous study of adapting and fine-tuning diffusion models using gradient guidance from an optimization perspective. Specifically, we first understand why naive gradient guidance does not lead to meaningful optimization performance. Built upon the insights gained from the analysis, we propose gradient guidance that is proven to preserve generated data structures and simultaneously achieve strong optimization guarantees on adapted variants of diffusion models.
We are aware of two recent works (Uehara et al., 2024; Marion et al., 2024) studying adapting the output distribution of diffusion models to a target reward function. In particular, they define a reward function with respect to the output distribution of the diffusion model. Given a pre-trained diffusion model, for instance, Uehara et al. (2024) utilizes a KL-divergence regularizer penalizing the deviation to the pre-trained model for preventing overfitting in fine-tuning. Through some explicit computation, Uehara et al. (2024); Marion et al. (2024) identify the proper guidance term to adapt the pre-trained model. Yet a sophisticated estimation procedure is needed to find the guidance term in Uehara et al. (2024); our gradient guidance enjoys much simplicity and efficacy, as we demonstrated in both theory and experiments.
3 Preliminaries: Diffusion Models
@@
Score-based diffusion models capture the distribution of pre-training data by learning a sequence of transformations to generate new samples from noise (Song et al., 2020c). A forward stochastic process progressively adds noise to data, whose sample trajectories are used to train the score function. To generate new samples, a backward denoising process starts from sampling pure noise and gradually transforms the noise guided by the learned score function.
Forward Process
@@
The forward process of diffusion models initializes with $X_{0}\in\mathbb{R}^{D}$, a random variable drawn from the pre-training data $\mathcal{D}$. It introduces noise to via an Ornstein-Uhlenbeck process, i.e.,
@@
Backward Process
If reversing the time of the forward process, we can reconstruct the original distribution of the data from pure noise. With $(\overline{W}_{t})_{t\geq 0}$ being another independent Wiener process, the backward SDE below (Anderson, 1982) reverses the time in the forward SDE (1),
@@
Score Matching
To learn the unknown score function $\nabla\log p_{t}(\cdot)$, it is common to train a score network $s_{\theta}(x,t)$ using samples generated by the forward process. Let $\mathcal{D}$ denote the data for training. Then the score network is learned by minimizing the following loss:
@@
where ${\mathcal{S}}$ is a given function class, ${\mathbb{E}}_{\mathcal{D}}$ denotes the empirical expectation over training data $\mathcal{D}$ and $\mathbb{E}_{x_{t}|x_{0}}$ denotes condition expectation over the forward process, $\phi_{t}(x_{t}|x_{0})$ is the Gaussian transition kernel, i.e., $\frac{1}{(2\pi h(t))^{D/2}}\exp(-\frac{\lVert x_{t}-\alpha(t)x_{0}\rVert^{2}}{2h(t)})$.
Generation and Guided Generation
Given a pre-trained score function $s_{\theta}$, one generates new samples by simulating the backward process (2) with the true score replaced by $s_{\theta}$. Further, one can add additional guidance to the backward SDE to steer its output distribution towards specific properties of interest. Module 1 formalizes the generation process and guided generation process using a pre-trained diffusion model.
@@
Conditional Generation
Suppose the goal is to generate $X$ with a desired property $Y=y$ from the distribution $P(X|Y=y)$. To this end, one needs the conditional score function $\nabla_{x_{t}}\log p_{t}(x_{t}\mid y)$, as a replacement of the unconditioned score $\nabla_{x_{t}}\log p_{t}(x_{t})$. The Bayes rule gives
@@
When a pre-trained score network $s_{\theta}(x_{t},t)\approx\nabla\log p_{t}(x_{t})$, the remaining task is to estimate $\nabla_{x_{t}}\log p_{t}(y\mid x_{t})$ and add it as a “guidance” G to the backward process (Module 1).
Classifier and Classifier-Free Guidance
Classifier guidance (Song et al., 2020c; Dhariwal and Nichol, 2021) is a approach for sampling from $P(X|Y=y)$ when $Y$ is a discrete label. This method estimates $\nabla_{x_{t}}\log p_{t}(y\mid x_{t})$ by training auxiliary classifiers, denoted as $\hat{p}(y\mid x_{t},t)$, and then computing the gradient of the classifier logits as the guidance, i.e., $\texttt{G}(x_{t},t)=\nabla_{x_{t}}\log\hat{p}(y\mid x_{t},t)$. An alternative is the classifier-free guidance method (Ho and Salimans, 2022), which jointly trains a conditional and an unconditional diffusion model, and combine the two score estimates via a form of guidance to generate samples.
Notations
For a random variable $X$, $P_{x}$ represents its distribution, and $p(x)$ denotes its density function. For $X$, $Y$ jointly distributed, $P(X\mid Y=y)$ denotes the conditional distribution, and $p(x\mid y)$ denotes its density function. We use the notation $\mathbb{E}[x\mid y]$ for the conditional expectation. Let $\mathcal{D}$ be the pre-training data, and let ${\mathbb{E}}_{\mathcal{D}}$ be the empirical expectation over $\mathcal{D}$. Let $\bar{\mu}$ and $\bar{\Sigma}$ denote the data’s empirical mean and covariance matrix, i.e., $\bar{\mu}:={\mathbb{E}}_{x\in\mathcal{D}}[x]$ and $\bar{\Sigma}:={\mathbb{E}}_{x\in\mathcal{D}}\left[(x-\bar{\mu})(x-\bar{\mu})^{\top}\right]$. For a matrix $A$, we denote by $\hbox{Span}(A)$ the subspace spanned by its column vectors. For a square matrix $A$, we denote by $A^{-1}$ its inverse or Moore–Penrose inverse. For any differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, $\nabla f\in\mathbb{R}^{m\times n}$ denotes Jacobian matrix, i.e., $(\nabla f)_{ij}=\frac{\partial f_{i}(x)}{\partial x_{j}}.$

### v2

@@ Abstract
Diffusion models have demonstrated empirical successes in various applications and can be adapted to task-specific needs via guidance. This paper studies a form of gradient guidance for adapting a pre-trained diffusion model towards optimizing user-specified objectives. We establish a mathematical framework for guided diffusion to systematically study its optimization theory and algorithmic design. Our theoretical analysis spots a strong link between guided diffusion models and optimization: gradient-guided diffusion models are essentially sampling solutions to a regularized optimization problem, where the regularization is imposed by the pre-training data. As for guidance design, directly bringing in the gradient of an external objective function as guidance would jeopardize the structure in generated samples. We investigate a modified form of gradient guidance based on a forward prediction loss, which leverages the information in pre-trained score functions and provably preserves the latent structure. We further consider an iteratively fine-tuned version of gradient-guided diffusion where guidance and score network are both updated with newly generated samples. This process mimics a first-order optimization iteration in expectation, for which we proved $\tilde{\mathcal{O}}(1/K)$ convergence rate to the global optimum when the objective function is concave. Our code will be released at https://github.com/yukang123/GGDMOptim.git.
@@ 1 Introduction
Diffusion models have emerged as a significant advancement in the field of generative artificial intelligence, offering state-of-the-art performance in image generation [55, 53, 21]. These models operate by gradually transforming a random noise into a structured output, utilizing the score function learned from data. One of the key advantages of diffusion models is their flexibility which allows controlled sample generation for task-specific interest, excelling diffusion models in a wide range of applications, such as content creation, sequential decision making, protein engineering [35, 1, 27].
Controlling the generation of large generative models stands at the forefront of AI. Guidance and fine-tuning are two most prevalent approaches for controlling the generation of diffusion models. Unlike fine-tuning which changes the weights of pre-trained models, guidance mechanism enables a more directed and flexible control. Adding gradient-based guidance during inference was pioneered by classifier guidance [57, 21], which involves training a time-dependent classifier. Diffusion Posterior Sampling (DPS) [15] introduced a fully training-free form of gradient-based guidance, which removes the dependence on time. This method has since been explored in numerous empirical studies [16, 34, 40, 62, 54, 67, 4]. However, despite these empirical successes, significant gaps remain in the theoretical understanding and guarantees of gradient-based guidance in diffusion models.
@@
Suppose we have a pre-trained diffusion model that can generate new samples faithfully, maintaining the latent structure of data. We study the problem of adapting this diffusion model to generate new samples that optimize task-specific objectives, while maintaining the learned structure in new samples. This problem has a strong connection to classic optimization, guided diffusion offers new possibilities to optimize complex design variables such as images, videos, proteins, and genomes [7, 63, 39] in a generative fashion. More comprehensive exposure to this middle ground can be found in recent surveys [66, 11, 28].
Given the optimization nature of this problem, it’s critical to answer the following theoretical questions from an optimization perspective: (i) Why doesn’t simply applying the gradient of the objective function w.r.t. the noised sample work? (ii) How to add a guidance signal to improve the target objective without compromising the sample quality? (iii) Can one guarantee the optimization properties of new samples generated by guided diffusion? (iv) What are the limits of adaptability in these guided models?
@@
Scope and Contribution.
In this paper, we investigate guided diffusion from an optimization perspective. To answer the four questions above, we propose an algorithmic framework, see Figure 1 for an illustration. Our main contributions are summarized as follows:
$\bullet$ Study structure-preserving guidance. We study the design of guidance under structural data distribution belonging to a latent low-dimensional subspace (1). We diagnose the failure of naive gradient guidance and study the theoretical aspects of guidance based on forward prediction loss (Definition 1), which provably preserves any low-dimensional subspace structure (Theorem 1).
$\bullet$ Establish a mathematical framework of guided diffusion. We build a mathematical framework for guided diffusion, which facilitates algorithm analysis and theory establishment. We propose and analyze an iterative guided diffusion using gradient queries on new samples (Algorithm 1; Figure 1 with fine-tuning block off). We give the first convergence theory showing generated samples converge to a regularized optimal solution w.r.t the objective (Theorem 4) with linear score class (12). The regularization is imposed by the pre-trained diffusion model, revealing a fundamental limit of adapting pre-trained diffusion models with guidance.
@@
$\bullet$ Provide rate-matching optimization theory. Furthermore, we propose an adaptive gradient-guided diffusion, where both pre-trained score network and guidance are iteratively updated using self-generated samples (Algorithm 2; Figure 1 with the fine-tuning block turned on). We show in expectation its iteration converges to a global optima within the latent subspace, at a rate of $\tilde{\mathcal{O}}(1/K)$ (Theorems 3, $K$ is ##\## of iterations), matching the classical convergence in convex optimization.
$\bullet$ Provide experimental justification. Simulation and image experiments are provided in Section 7 to support theoretical findings on latent structure-preserving and optimization convergence.
2 Related Works
To summarize the related work, we first give an overview of empirical studies relevant to our objective. We then discuss the theory of diffusion models, to which our main contribution is focused. Other related topics, such as direct latent optimization in diffusion models and a detailed review of sampling and statistical theory of diffusion models, are deferred to Appendix A.
Classifier Guidance and Training-free Guidance.
[21] introduced classifier-based guidance, steering pre-trained diffusion models towards a particular class during inference. This method offers flexibility by avoiding task-specific fine-tuning, but still requires training a time-dependent classifier. Training-free guidance methods [15, 54, 67, 4, 30, 47, 26] eliminate the need for a time-dependent classifier, using only off-the-shelf loss guidance during inference. [15, 54, 30, 26] is a line of works solving inverse problems on image and [67, 47] aims for guided/conditional image generation. Though not originally developed for solving optimization problems, [15, 67] both propose a similar guidance to ours: taking gradient on the predicted clean data $x_{0}$ with respect to corrupted $x_{t}$. Differently, our paper presents the first rigorous theoretical analysis of this gradient-based guidance approach. Furthermore, we propose an algorithm that iteratively applies the guidance as a module to the local linearization of the optimization objective, demonstrating provable convergence guarantees.
Fine-tuning of Diffusion Models.
Several methods for fine-tuning diffusion models to optimize downstream reward functions include RL-based fine-tuning [7, 24] and direct backpropagation to rewards [17, 49, 65, 60]. However, these approaches often suffer from high computational costs and catastrophic forgetting in pre-trained models. Our guidance method is training-free and applied during the inference phase, eliminating the need to fine-tune diffusion models.
Theory of Diffusion Models.
Current theory works primarily focus on unconditional diffusion models. Several studies demonstrate that the distributions generated by diffusion models closely approximate the true data distribution, provided the score function is accurately estimated [19, 2, 8, 36, 12, 37, 14, 13, 6]. For conditional diffusion models, [68, 25] establish sample complexity bounds for learning generic conditional distributions. Our novel analysis establishes a connection between the sampling process in gradient-based guided diffusion and a proximal gradient step, providing convergence guarantees.
3 Preliminaries: Diffusion Models and Guidance
@@
Score-based diffusion models capture the distribution of pre-training data by learning a sequence of transformations to generate new samples from noise [57]. A diffusion model comprises a forward and a backward process, for which we give a review as follows.
Forward Process.
@@
The forward process progressively adds noise to data, and then the sample trajectories are used to train the score function. The forward process initializes with $X_{0}\in\mathbb{R}^{D}$, a random variable drawn from the pre-training data $\mathcal{D}$. It introduces noise via an Ornstein-Uhlenbeck process, i.e.,
@@
Backward Process.
If reversing the time of the forward process, we can reconstruct the original distribution of the data from pure noise. With $(\overline{W}_{t})_{t\geq 0}$ being another independent Wiener process, the backward SDE below [3] reverses the time in the forward SDE (1),
@@
Score Matching.
To learn the unknown score function $\nabla\log p_{t}(\cdot)$, we train a score network $s_{\theta}(x,t)$ using samples from forward process. Let $\mathcal{D}$ denote the data for training. Then the score network is learned by minimizing the following loss:
@@
where ${\mathcal{S}}$ is a given function class, ${\mathbb{E}}_{\mathcal{D}}$ denotes the empirical expectation over training data $\mathcal{D}$ and $\mathbb{E}_{x_{t}|x_{0}}$ denotes condition expectation over the forward process, $\phi_{t}(x_{t}|x_{0})$ is the Gaussian transition kernel, i.e., $(2\pi h(t))^{-D/2}\exp(-{\lVert x_{t}-\alpha(t)x_{0}\rVert^{2}}/\left(2h(t)\right))$.
Generation and Guided Generation.
Given a pre-trained score function $s_{\theta}$, one generates samples by the backward process (2) with the true score replaced by $s_{\theta}$. Further, one can add additional guidance to steer its output distribution towards specific properties, as formulated in Module 1.
@@
A common goal of guided generation (Module 1) is to generate $X$ with a desired property $Y=y$ from the distribution $P(X|Y=y)$. To this end, it essentially needs to learn the conditional score function $\nabla_{x_{t}}\log p_{t}(x_{t}\mid y)$. The Bayes rule gives
@@
When a pre-trained score network $s_{\theta}(x_{t},t)\approx\nabla\log p_{t}(x_{t})$, what remains is to estimate $\nabla_{x_{t}}\log p_{t}(y\mid x_{t})$ and add it as the guidance term G to the backward process in Module 1.
Classifier and Classifier-Free Guidance.
Classifier guidance [57, 21] samples from $P(X|Y=y)$ when $Y$ is a discrete label. This method estimates $\nabla_{x_{t}}\log p_{t}(y\mid x_{t})$ by training auxiliary classifiers, denoted as $\hat{p}(y\mid x_{t},t)$, and then computing the gradient of the classifier logits as the guidance, i.e., $\texttt{G}(x_{t},t)=\nabla_{x_{t}}\log\hat{p}(y\mid x_{t},t)$. Alternatively, classifier-free guidance [31] jointly trains a conditional and an unconditional diffusion model, combining their score estimates to generate samples.
Notations.
For a random variable $X$, $P_{x}$ represents its distribution, and $p(x)$ denotes its density. For $X$, $Y$ jointly distributed, $P(X\mid Y=y)$ denotes the conditional distribution, and $p(x\mid y)$ its density. The conditional expectation is denoted as $\mathbb{E}[x\mid y]$. Let $\mathcal{D}$ be the pre-training data, and let ${\mathbb{E}}_{\mathcal{D}}$ be the empirical expectation over $\mathcal{D}$. The empirical mean and covariance matrix are $\bar{\mu}:={\mathbb{E}}_{x\in\mathcal{D}}[x]$ and $\bar{\Sigma}:={\mathbb{E}}_{x\in\mathcal{D}}[(x-\bar{\mu})(x-\bar{\mu})^{\top}]$. For a matrix $A$, $\operatorname{Span}(A)$ denotes the subspace spanned by its column vectors, and for a square matrix $A$, $A^{-1}$ denotes its inverse or Moore–Penrose inverse. For any differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, $\nabla f\in\mathbb{R}^{m\times n}$ denotes Jacobian matrix, i.e., $(\nabla f)_{ij}=\frac{\partial f_{i}(x)}{\partial x_{j}}.$